# Can Quasi Periodic Oscillations Encode Traces of Black Hole Phase Transitions ?

### v1

@@ Abstract
Finding observational evidence or imprints of black hole phase transitions is a promising and active area of research. Aiming to contribute to this direction, in this work, we probe the well-known thermodynamic phase structure of the Reissner Nordström Anti de Sitter (RN-AdS) black hole through the lens of its quasi-periodic oscillations (QPOs). Can QPOs be influenced by black hole phase transitions? Do they carry any signature of such transitions in their observational patterns? These were the central questions guiding our study. By analyzing the upper and lower QPO frequencies within the Relativistic Precession (RP) model, we observe that the frequency-temperature relationship reflects distinguishing features corresponding to the small, intermediate, and large black hole phases, offering insights into their stability properties. The analysis is further extended to other QPO models, including the Warped Disk (WD) and Epicyclic Resonance (ER) models. The presence of distinct phase structures remains evident across all models which is an encouraging outcome that may be worth exploring further in the context of other black hole systems.
@@ II QPO and phase transition
II.1 Massless particle (null geodesic) : RN black hole case
@@
From Eq. (13) it is evident that $r_{0}$ is independent of L Using the lapse function of RN AdS black hole, Eq.(13) is solved for $r_{0}$ Now substituting the expression of mass in terms of event horizon radius $r_{+}$, we obtained the following relation between radius of the unstable circular orbit $r_{0}$ and event horizon radius $r_{+}$ as
@@
II.2 QPO Models
In this section, we explore the behavior of twin-peak quasi-periodic oscillations (QPOs) in the background of Reissner–Nordström (RN) black holes. The upper ($\nu_{U}$) and lower ($\nu_{L}$) QPO frequencies are modeled as functions of the radial coordinate and black hole parameters, guided by several well-established theoretical frameworks [78].
@@ II.2 QPO Models
II.3 Signatures of Phase transition
@@
As illustrated in Fig. 4, the solid red curve represents the SBH branch. In this regime, both $\nu_{U}$ and $\nu_{L}$ decrease with rising temperature, indicating that the frequency of a particle orbiting the black hole reduces as the system heats up. This monotonic decrease is a hallmark of stability in the small black hole phase.
In contrast, along the IBH branch(solid blue curve), both frequencies increase with temperature, suggesting an unstable regime where the particle experiences stronger oscillations — a behavior consistent with thermodynamic instability. Finally, in the LBH branch (solid black curve) the frequencies begin to decrease once more and eventually saturate, reflecting a return to stability in the large black hole phase. When the black hole charge exceeds its critical value, the system exhibits a single thermodynamically stable phase across the entire temperature domain. This is reflected in the frequency behavior as well: for $Q$ above the critical threshold, both $\nu_{U}$ and $\nu_{L}$ decrease monotonically with temperature and asymptotically approach constant values. This smooth decline further confirms the absence of phase transitions and supports the thermodynamic stability of the system in this regime.
This analysis highlights that the QPO behavior of a test particle can serve as an effective probe of black hole phase transitions and their stability. The connection between dynamical frequencies and thermodynamic stability suggests a compelling observational signature for identifying black hole phases. However, further observational data will be crucial to solidify this correspondence.
To reinforce our findings, we provide a numerical table below summarizing the QPO frequencies across the three branches, which further illustrates the thermodynamic phase structure inherent in RN AdS black holes.
Black Hole Phase Temperature ($T$) Upper Frequency ($\nu_{U}$) [Hz] Lower Frequency ($\nu_{L}$) [Hz]
$Q<Q_{c}$ (Three-Phase Structure)
0.0119366 164709 47810.3
0.41958 159488 41240.9
SBH 0.573522 149007 30901.6
0.625255 137596 22577.1
0.633382 $(T_{C})$ 129336 17886.8
0.632781 126880 16665.2
0.620704 117308 12539.0
IBH 0.420766 66966.8 1710.48
0.329510 49723.0 514.412
0.274613 $(T_{C})$ 36308.2 70.2909
0.274948 35784.1 59.5509
0.318111 32972.8 9.14767
LBH 0.742716 32382.3 0.0399761
1.209580 32379.8 0.00212883
1.682490 32379.6 0.000293692
$Q>Q_{c}$ (Single Stable Phase)
0.119366 34981.8 5588.86
0.156414 34701.6 4759.03
Stable BH 0.183765 34399.5 3948.53
0.205161 34108.0 3235.86
0.222794 33843.0 2639.13
Table 1: Numerical values of upper and lower QPO frequencies of massless particle across different black hole phases for $Q<Q_{c}$ (SBH, IBH, LBH) and $Q>Q_{c}$ (single-phase stable branch). The trends in QPO behavior reflect the underlying thermodynamic phase structure and stability patterns of RN-AdS black holes.
We have also performed our analysis using the ER2, ER3, ER4 models as well as the Warped Disk (WD) model. All these QPO models exhibit consistent qualitative behavior and successfully capture the features of black hole phase transitions and the associated stability structure. For the sake of brevity, we omit the detailed results here. It is worth noting, however, that while the overall trends remain coherent across models, the actual frequency values may vary. Among them, the ER2 model yields the highest frequency values. Across all five models, the behavior of the upper and lower QPO frequencies as functions of temperature remains qualitatively similar: in the SBH phase, frequencies decrease with temperature, indicating stable motion; in the IBH phase, the frequencies rise, signaling instability; and in the LBH phase, they gradually level off and approach constant values in the asymptotic limit, once again reflecting stability.
III Massive particle(timelike geodesic)
@@
In the SBH region, both $\nu_{U}$ and $\nu_{L}$ decrease with increasing temperature, indicating dynamical stability. In the IBH branch, the frequencies increase with temperature, which corresponds to an unstable regime where orbital motion becomes more sensitive to perturbations. Interestingly, as the system transitions into the LBH branch, the lower frequency $\nu_{L}$ for massive particles drops sharply and reaches zero exactly at the onset of the LBH phase. This vanishing of $\nu_{L}$ signifies that the particle enters a state of ultimate stability, where radial oscillations effectively cease. Beyond this point, $\nu_{L}$ becomes negative, which is unphysical and indicates the breakdown of the oscillatory interpretation in this regime. In contrast, the upper frequency $\nu_{U}$ continues to follow the same trend as in the massless case, decreasing monotonically and asymptotically approaching a constant value.
@@
For charge values above the critical threshold ($Q>Q_{c}$), the temperature-frequency relation is smooth and monotonic, with no indication of a phase transition. In this regime, both frequencies evolve predictably: $\nu_{U}$ decreases gradually with temperature, while $\nu_{L}$ remains positive and asymptotically approaches zero from above, indicating a stable configuration throughout.
In summary, while the overall qualitative behavior of QPO frequencies for massive particles is consistent with the massless case, the vanishing of $\nu_{L}$ at the LBH threshold introduces a distinctive physical signature. A numerical summary of the results for massive particles is provided in the table below.
Black Hole Phases Temperature (T) Lower Frequency ($\nu_{L}$) [Hz] Upper Frequency ($\nu_{U}$) [Hz]
0.0119366 47810.1 164711.
0.41958 41240.4 159490.
SBH 0.573522 30900.7 149009.
0.625255 22575.7 137599.
0.633382($T_{C}$) 17885.1 129339.
0.632781 16663.4 126883.
0.620704 12536.7 117311.
IBH 0.420766 1701.11 66975.1
0.32951 486.214 49738.1
0.280928 0 39376.7
0.274613($T_{C}$) -338.359 36366.1
$Q>Q_{c}$ (Single Stable Phase)
0.119366 4966.37 35053.9
0.156414 3946.51 34783.2
Stable Phase 0.183765 2845.78 34494.1
0.205161 1715.0 34218.4
0.222794 537.2 33972.2
Table 2: Numerical values of the temperature of RN AdS black hole and QPO frequencies for massive particle
We have also extended our analysis to other QPO models, including the Warped Disk (WD) model and the Epicyclic Resonance (ER) models. Among them, the WD model exhibits behavior that closely mirrors that of the Relativistic Precession (RP) model. Both models consistently capture the thermodynamic stability pattern of black hole phases and support our interpretations based on frequency evolution. In contrast, the ER modelsnamely ER2, ER3, and ER4 do not fully align with the expected behavior. Specifically, in all ER models, the frequencies associated with the large black hole (LBH) phase exhibit an increasing trend with temperature, which contradicts the physical interpretation drawn from the RP and WD models. While the presence of distinct phase structures remains evident across all models, the discrepancies observed in the ER frameworks raise an interesting question that needs to be addressed in our future works.
@@ IV Summary and Concluding Remarks
In this work, we probe the well-known thermodynamic phase structure of the RN-AdS black hole through the lens of its quasi-periodic oscillations (QPOs). Can QPOs be influenced by black hole phase transitions? Do they carry any signature of such transitions in their observational pattern? These were the central questions guiding our study. Beginning with the Relativistic Precession (RP) model, we demonstrated that the behavior of both upper and lower QPO frequencies, when plotted against temperature reveals a clear imprint of the underlying phase structure for subcritical charge values. Specifically, the frequencies trace out the small, intermediate, and large black hole branches in a manner that reflects their respective thermodynamic stability. In the small and large black hole phases, the frequencies decrease with temperature, indicating stability, whereas the intermediate branch is characterized by increasing frequencies, signaling instability. For charge values above the critical point, our results show a smooth, monotonic behavior of the frequencies with no evidence of phase transitions, further supporting the interpretation of a single thermodynamically stable phase in this regime. We have also extended our analysis to massive particles and confirmed that the qualitative features of the frequency-temperature relationship remain intact. In addition to the RP model, we examined the behavior of QPO frequencies under the Warped Disk (WD) and Epicyclic Resonance (ER) models. For massless particles, all models including ER2, ER3, ER4, and WD exhibit consistent qualitative trends. In the massive particle case, however, the ER models exhibit noticeable deviations most prominently in the LBH branch, where the QPO frequencies increase with temperature, in sharp contrast to the stability characteristics captured by the RP and WD models. On the other hand, the WD model continues to mirror the RP model and successfully reflects the expected stability pattern. Although the presence of distinct phase structures remains evident across all models which is an encouraging feature for further exploration in other black hole systems. The precise relationship between thermodynamic stability and QPO behavior may vary depending on the black hole geometry or the choice of thermodynamic ensemble. This remains an open question, inviting deeper investigation in future studies.

### v2

# Probing Black Hole Phase Transitions through Quasi-Periodic Oscillations

@@ Abstract
In this work, we probe the well known thermodynamic phase structure of black hole through the lens of its quasi-periodic oscillations (QPOs). Can QPOs be influenced by black hole phase transitions? Do they carry any signature of such transitions in their observational patterns? These were the central questions guiding our study. Using both RN AdS and Kerr black hole backgrounds across different QPO models, we analyzed the behavior of upper and lower QPO frequencies as functions of the Hawking temperature. Our results shows that QPO frequencies trace out distinct thermodynamic phases and also reflect their stability properties. As the black hole transitions between different thermodynamic phases, the trend of QPO frequencies with respect to temperature also shifts. Due to lack of of observational data, the present work is primarily more on the mathematical side, as the underlying mechanism responsible for the Hawking temperature has not yet been fully understood or experimentally verified. Moreover, given the speculative nature of black hole phase transitions, it would be unfair to claim that our results establish a definitive connection between an observable quantity such as the QPO frequency and the thermodynamic phase behavior of black holes. Nevertheless, our analysis suggests a possibility that changes in black hole geometry could be one of the contributing factors influencing QPO behavior.
@@ I Introduction
The objective of this work is to explore possible imprints of black hole thermodynamic properties using quasi-periodic oscillations (QPOs). As a standard practice we start with the Reissner-Nordström Anti-de Sitter (RN-AdS) black hole. Although the RN-AdS black hole is not directly representative of astrophysical black holes due to its static nature and asymptotically non-flat boundary but it serves as an excellent theoretical model for exploring black hole phase transitions. In particular, the RN-AdS geometry possesses a well-understood phase structure, including Van der Waals-like transitions and clear thermodynamic branches, which makes it an ideal testbed for probing whether quasi-periodic oscillation (QPO) frequencies can encode information about underlying black hole phase transitions. The aim of this study is thus conceptual : to investigate if QPO signatures can reflect thermodynamic phenomena in black hole spacetimes.The background geometry will be upgraded to the Kerr spacetime later in this work. Since Kerr black holes are rotating and asymptotically flat, they more accurately describe observed black hole systems, particularly those involving accretion and QPO generation. It is important to mention that, in this work, we focus solely on the relationship between Hawking temperature and QPO frequencies, treating Hawking temperature as a theoretical control parameter reflecting the black hole’s thermodynamic state. We do not incorporate the effects of the temperature of accretion disk in our analysis.
II RN $AdS$ black hole
II.1 Thermodynamics of RN AdS black hole
@@ II.2 QPO and phase transition
II.3 Massless particle (null geodesic) : RN black hole case
@@
From Eq. (13) it is evident that $r_{0}$ is independent of L Using the lapse function of RN AdS black hole, Eq.(13) is solved for $r_{0}$ . Now substituting the expression of mass in terms of event horizon radius $r_{+}$, we obtained the following relation between radius of the unstable circular orbit $r_{0}$ and event horizon radius $r_{+}$ as
@@
II.4 QPO Models
In this section, we explore the behavior of twin-peak quasi-periodic oscillations (QPOs) in the background of RN AdS black holes. The upper ($\nu_{U}$) and lower ($\nu_{L}$) QPO frequencies are modeled as functions of the radial coordinate and black hole parameters, guided by several well-established theoretical frameworks [78].
@@ II.4 QPO Models
II.5 Signatures of Phase transition
@@
As illustrated in Fig. 5, the solid red curve represents the SBH branch. In this regime, both $\nu_{U}$ and $\nu_{L}$ decrease with rising temperature, indicating that the frequency of a particle orbiting the black hole reduces as the system heats up. In contrast, along the IBH branch(solid blue curve), both frequencies increase with temperature, suggesting an unstable regime where the particle experiences stronger oscillations. Finally, in the LBH branch (solid black curve) the frequencies begin to decrease once more and eventually saturate. When the black hole charge exceeds its critical value, the system exhibits a single thermodynamically stable phase across the entire temperature domain. This is reflected in the frequency behavior as well: for $Q$ above the critical threshold, both $\nu_{U}$ and $\nu_{L}$ decrease monotonically with temperature and asymptotically approach constant values.
It is evident that the QPO behavior of a test particle can serve as an effective probe of black hole phase transitions and their stability. The connection between dynamical frequencies and thermodynamic stability suggests a compelling observational signature for identifying black hole phases. However, further observational data will be crucial to solidify this correspondence.
II.6 Massive particle(timelike geodesic)
@@
In the SBH and LBH region, both $\nu_{U}$ and $\nu_{L}$ decrease with increasing temperature. In the IBH branch, the frequencies increase with temperature. For charge values above the critical threshold ($Q>Q_{c}$), the temperature-frequency relation is smooth and monotonic, with no indication of a phase transition. In this regime, both frequencies evolve predictably: $\nu_{U}$ decreases gradually with temperature, while $\nu_{L}$ remains positive and asymptotically approaches zero from above. In summary, the overall qualitative behavior of QPO frequencies for massive particles is consistent with the massless case.
It is important to note that the QPO frequencies observed are believed to be influenced by the temperature and dynamics of the accretion disk, which can reach values of $10^{6}$$10^{7}$ K and is accessible via X-ray spectral observations. In contrast, the Hawking temperature of a black hole, is an extremely small quantity and has no direct observational signature as of now. However, changes in the Hawking temperature are associated with changes in the geometry of the black hole spacetime, particularly near the event horizon. Since QPO frequencies are sensitive to the spacetime structure in the inner accretion region, these geometric changes can influence their behavior. We do not suggest a direct physical connection between the two, but rather use the temperature as a thermodynamic indicator of the black hole’s geometric phase. By plotting QPO frequencies directly against Hawking temperature, we aim to isolate the influence of the black hole geometry on the QPO behavior. Although the exact physical mechanism connecting the two remains complex and not fully understood, the way QPO frequencies change across different thermodynamic phases reveals meaningful patterns. These results suggest that QPOs could carry observable signatures of black hole phase transitions.
We have also performed our analysis using the ER2, ER3, ER4 models as well as the Warped Disk (WD) model for both mass and massless particle case. All these QPO models exhibit consistent qualitative behavior and successfully capture the features of black hole phase transitions and the associated stability structure. For the sake of brevity, we omit the detailed results here while these analysis is shown for Kerr black hole as it will be more physically meaningful. It is worth noting, however, that while the overall trends remain coherent across models, the actual frequency values may vary. Across all five models, the behavior of the upper and lower QPO frequencies as functions of temperature remains qualitatively similar: in the SBH phase, frequencies decrease with temperature, indicating stable motion; in the IBH phase, the frequencies rise, signaling instability; and in the LBH phase, they gradually level off and approach constant values in the asymptotic limit.
III Kerr black hole
Figure 5: Behavior of the quasi-periodic oscillation (QPO) frequencies of both upper $\nu_{U}$ and lower $\nu_{L}$ for Kerr black holes as functions of temperature, across different black hole branches in RP model : small (SBH) and large (LBH). Here, we have considered $a=0.5$.
To ensure astrophysical relevance, we adopt the Kerr black hole geometry as the background spacetime. We have considered the black holes to be rotating and asymptotically flat, making them more consistent with the nature of astrophysical black holes observed in X-ray binaries and active galactic nuclei. Rotation is a crucial feature in accreting black hole systems, as the presence of an accretion disk inherently implies angular momentum. In this case we only consider the massive particle case, as it will be more physically meaningful.
III.1 Thermodynamics of Kerr Black hole
The mass of the black hole is
$M=\frac{a^{2}+r_{+}^{2}}{2r_{+}}$ (24)
The temperature is calculated as
$T=\frac{r_{+}^{2}-a^{2}}{4\pi r_{+}^{3}}$ (25)
Figure 6: Temperature of the Kerr black hole with respect to event horizon radius
In the temperature plot, we observe two black hole branch. As the rotation $a$ value increases, the temperature value decreases. The SBH branch is found to be thermodynamically stable while the LBH branch is found to be unstable. We next study the QPO profiles.
The effective potential in case of a massive particle orbiting around a Kerr black hole is
$V_{eff}=\\ \frac{-\left(\text{E}\left(a^{2}+r^{2}\right)-aL\Xi\right)^{2}+\Delta(r)(a\text{E}-L\Xi)^{2}+r^{2}(-\Delta(r))}{r^{4}}$ (26)
Accordingly, the squared radial and vertical frequencies in case of Kerr black holes can be calculated as
$\Omega_{r}^{2}=\frac{Mr\left(r(r-6M)-3a^{2}\right)+8a(Mr)^{3/2}}{r^{2}\left(a\sqrt{Mr}+r^{2}\right)^{2}}$ (27)
$\Omega_{\theta}^{2}=\frac{3a^{2}Mr-4aMr\sqrt{Mr}+Mr^{3}}{\left(ar\sqrt{Mr}+r^{3}\right)^{2}}$ (28)
$\Omega_{\phi}^{2}=\frac{Mr}{\left(a\sqrt{Mr}+r^{2}\right)^{2}}$ (29)
Next, we revisit the Relativistic Precession (RP) model to compute the upper and lower QPO frequencies as functions of the event horizon radius $r_{+}$. Using these results, we construct a parametric plot between the black hole temperature and the corresponding QPO frequencies. The plot reveals the existence of two distinct black hole branches. For the small black hole (SBH) branch, we observe that the QPO frequencies decrease with increasing temperature. This monotonic behavior indicates thermodynamic stability, as the system responds smoothly to changes in temperature. In contrast, along the large black hole (LBH) branch, the QPO frequencies exhibit an increasing trend with temperature. This reversal in behavior suggests thermodynamic instability in the LBH regime.
Our analysis reveals that the slope of the QPO frequency–temperature relation, as represented by $\arctan(\nu_{L}/T)$, is notably influenced by the black hole’s spin parameter $a$. As the rotation increases, we observe that the slope steepens in specific regions, especially near critical points where frequency behavior changes rapidly. A steeper slope indicates a stronger sensitivity of the QPO frequency to variations in temperature, suggesting that even small thermodynamic changes in the black hole result in noticeable shifts in QPO characteristics. Moreover, increasing spin appears to shift the onset of these rapid variations to larger horizon radii, implying a delay in the transition behavior.
The signatures observed in the QPO frequency–temperature relations likely arise from an indirect but interesting connection between black hole thermodynamics and the motion of matter in the surrounding spacetime. When a black hole undergoes a thermodynamic transition such as moving from a small to a large black hole phase it changes the geometry near the event horizon. These geometric changes affect the effective potentials that determine how particles and fluid elements move around the black hole. Since QPO frequencies are directly related to this motion, any shift in the spacetime geometry can influence the observed oscillation frequencies. The black hole’s spin plays a key role in this process. Increasing the spin not only changes the temperature and thermodynamic behavior of the black hole, but also alters the shape of the spacetime and the stability of particle orbits. As a result, the response of QPO frequencies to changes in temperature becomes more sensitive in rapidly rotating black holes.
We performed our analysis across five distinct QPO models: the Relativistic Precession (RP) model, three variants of the Epicyclic Resonance (ER) model namely ER2, ER3, and ER4 and the Warped Disk (WD) model. For each model, we computed the upper and lower QPO frequencies as functions of the event horizon radius and black hole spin. Despite differences in their construction and frequency relations, all five models exhibit consistent qualitative behavior with respect to black hole thermodynamic structure as shown in Fig.7 In particular, the characteristic signatures of the small and large black hole branches, as well as the influence of rotation on QPO–temperature response, remain robust across models. Among these, the ER3 model yields the highest frequency values, as evident from Fig. 8, which aligns with its formulation involving higher-order resonances.
Refer to caption    Refer to caption
Refer to caption   Refer to caption
Figure 7: Behavior of the quasi-periodic oscillation (QPO) frequencies of both upper $\nu_{U}$ and lower $\nu_{L}$ for Kerr black holes as functions of temperature, across different QPO models. Here, we have considered $a=0.5$.
@@ IV Summary and Concluding Remarks
In this work, we treated the QPO frequencies as a way to probe the thermodynamic phase structure of the black hole. Can QPOs be influenced by black hole phase transitions? Do they carry any signature of such transitions in their observational pattern? These were the central questions guiding our study. We demonstrated that the behavior of both upper and lower QPO frequencies, when plotted against temperature reveals a clear imprint of the underlying phase structure. Specifically, the frequencies trace out different black hole branches in a manner that reflects their respective thermodynamic stability.
We considered two different black hole system for our analysis, the RN black hole and the Kerr black hole. Although the RN black hole is not astrophysically realistic, it served as a useful toy model with a well-understood thermodynamic phase structure, providing a standard starting point for investigating black hole phase transitions. We aimed to isolate how geometric changes associated with thermodynamic transitions affect the QPO behavior. Our results showed consistent patterns across different thermodynamic phases, suggesting that QPOs may carry indirect signatures of black hole phase transitions. This appears to arise from the way geometric changes near the event horizon effects the particle motion. For the Kerr black hole, we found that black hole spin significantly influenced the QPO–temperature relation, making the QPO frequencies more sensitive to thermodynamic changes in high-spin regimes. From the above study may be we can summarize as : if the slope of the QPO–temperature curve is negative or zero, the thermodynamic phase is stable whereas if the slope is positive, the phase becomes unstable. Mathematically, this condition can be expressed as:
$\begin{cases}\dfrac{d\Omega_{i}}{dT}\leq 0,&\text{Stable phase},\\[10.0pt] \dfrac{d\Omega_{i}}{dT}>0,&\text{Unstable phase}.\end{cases}$ (30)
Here $i$ refers to the upper or lower QPO frequencies. To establish the above relation more rigorously, a broader range of black hole systems should be examined. With a refined mathematical framework, it may be possible to extract more information about the nature of this relation and the underlying physical mechanism. A detailed investigation in this direction is reserved for future work.
It should be noted that the present findings may be purely mathematical in nature and may not possess direct physical relevance, as the Hawking temperature is not yet an observable quantity. However, if a physical correspondence does exist, then QPOs could serve as an effective tool to probe the Hawking temperature and the associated black hole phase transitions. This may further suggest that QPOs are influenced not only by accretion-related processes but also by the underlying thermodynamic state of the black hole, which governs its spacetime geometry.
QPOs observed in X-ray binaries have sparked a great deal of interest, and over the years, several models have been proposed to explain their origin. One of the leading ideas is the relativistic precession (RP) model, which links QPOs to the precession of matter orbiting in the strong gravitational field near the black hole. Another class of models, known as epicyclic resonance (ER) models, suggests that QPOs arise from resonances between different oscillation modes, specifically radial and vertical epicyclic motions, in the accretion disk. There are also disk oscillation models that attribute QPOs to global vibrations of the disk. Other interpretations involve warped disk dynamics, magnetic field instabilities, and beat-frequency mechanisms. While each model offers valuable insights, the true origin of QPOs remains uncertain, and it is likely that different mechanisms may be at play depending on the physical environment. In addition to these interpretations, our work suggests that the spacetime geometry of the black hole itself may also play a role in shaping QPO behavior. Changes in the Hawking temperature are closely tied to changes in the black hole’s geometry, particularly in the near-horizon region. Since QPO frequencies are highly sensitive to the structure of spacetime in the inner accretion zone, these geometric variations can significantly influence their behavior. In our interpretation, QPOs are affected not only by accretion-related processes but also by the underlying thermodynamic state of the black hole, which governs its geometry. This is clearly reflected in our results: as the black hole transitions between different thermodynamic phases, the trend of QPO frequencies with respect to temperature also shifts. These findings suggest that QPOs may carry imprints of the evolving geometry, and our work points to the possibility that changes in black hole geometry could be one of the contributing factors influencing QPO behavior.